OpenStudy (anonymous):

write in sigma notation 1/1 + 1/4 +1/9 +...1/100

OpenStudy (kc_kennylau):

\[\Large\sum_{i=1}^{10}\frac1{i^2}\]

OpenStudy (kc_kennylau):

When i=1, 1/i^2 = 1/1 When i=10, 1/i^2 = 1/100

OpenStudy (anonymous):

why is it i? o.o

OpenStudy (anonymous):

THANK YOU SO MUCH :D

OpenStudy (kc_kennylau):

You can use any variable

OpenStudy (anonymous):

ohh lol xD

OpenStudy (kc_kennylau):

no problem :)